Solve by the Addition Method.Indicate whether it is independent, inconsistent, or dependent.
1/3x - 1/6y = 1/3
1/6x + 1/4y =0

Answer by stanbon(75887)

Solve by the Addition Method.Indicate whether it is independent, inconsistent, or dependent.
1/3x - 1/6y = 1/3
1/6x + 1/4y =0
-----------------
Multiply thru the 2nd equation by 2 to get:
1st: (1/3)x - (1/6)y = (1/3)
2nd: (1/3)x + (1/2)y = 0
-------------------------
Subtract 2nd from 1st to get:
[(1/6)-(1/2)]y = (1/3)
(-1/3)y = 1/3
y = -1
--------------
Substitute to solve for "x":
(1/6)x = (-1/4)
x = (-1/4)*6 = -3/2
-------------------------
Solution: x = -3/2 ; y = 1
===================================
Since the lines meet in one point, the system is independent.
===================================
Cheers,
Stan H.